OpenStudy (anonymous):

x+24 divided by x+4 is greater than or equal to 4

OpenStudy (phi):

\[ \frac{x+24}{x+4} ≥ 4 \]

OpenStudy (phi):

you could add -4 to both sides \[ \frac{x+24}{x+4} -4 ≥ 0\\\frac{x+24}{x+4}+\frac{-4(x+4)}{x+4}≥ 0\\ \frac{x+24-4x-16}{x+4} ≥ 0 \\ \frac{-3x+8}{x+4} ≥ 0\ \]

OpenStudy (phi):

that fraction will be positive (that is, bigger than 0) when both the top and bottom are negative, or both are positive so you should test where -3x+8>0 and x+4>0 and where -3x +8< 0 and x+4<0